# How Far Can 100 Samples Go? Unlocking Overall Zero-Shot Multilingual Translation via Tiny Multi-Parallel Data

### 22 Jan 2024 v1

@@ 3.4 More Data or More Directions?
In Section 3.3, we showed that fine-tuning an English-centric model with bitext derived from NTREX (covering all directions) yields substantial overall zero-shot improvements. A natural assumption is that these improvements are triggered by the corresponding directional data. In this section, we investigate whether it is true, i.e., what will happen if we only cover a subset of language directions during fine-tuning.
We conduct experiments on the same English-centric model trained on EC30 (see Section 3.3) and control the scale of the fine-tuning data in the following two settings:
  1. (a)
    We conducted a random sampling of 100 multi-parallel NTREX sentences to construct pairwise data to cover all 870 directions [^5] These directions include English-centric ones while excluding Occitan-related ones as it is not available in NTREX.. Then, we randomly sampled {10%, 20%, 40%, 80%} directions for fine-tuning.
  2. (b)
    We fixed the 10% of directions as mentioned in (a) and conducted a random sampling of {100, 200, 400, 800} multi-parallel NTREX instances to construct the fine-tuning set for the corresponding directions.
@@
Figure 4: ChrF improvements of the upper bound and boosted models over the English-centric baseline on the Europarl-8 dataset. It is easy to see that the non-English capability of the boosted model is quite close to the upper bound (complete translation), meanwhile, it also holds the performance in English-centric directions.
We show all of the corresponding fine-tuning results in Figure 2. Surprisingly, when fixing the size of the multi-parallel samples as 100 and then increasing the fine-tuning directions from 10% to 80%, no improvement is observed for any resource group (Figure 2-a). Fine-tuning in randomly sampled 10% directions using 100 samples achieves comparable overall results to fine-tuning in all directions (Boost-100). However, when we fix the directions as 10%, and extend the multi-parallel sample size from 100 to 800, consistent improvements for all groups can be observed (Figure 2-b). It reveals that the overall performance is not sensitive to the count of directions, at least when the directional data extend to a certain scale, like 10%.
@@
3.5 How Close to the Upper Bound?
In this section, we empirically show how close to the upper bound the fine-tuning method can arrive. Here, we treat the performance of complete translation, i.e., training with fully multi-parallel data, as the upper bound, where identical scales of non-English pairwise data cover all of the directions that the English-centric counterparts can not cover.
We conduct experiments on Europarl-8 (see Section 3.1), where 8-way aligned data are available. We train English-centric and complete translation models based on it, respectively. The non-English performance of the complete translation model is seen as the "upper-bound" capability here where the information from the bitext in all directions is identical. Also, we report the results of fine-tuning the English-centric model with pairwise data constructed from NTREX that covers all directions.
In Figure 4, we show that 1) the upper-bound performance, i.e., that of the complete model, surpasses the baseline by a big margin, resulting in +17.4 average ChrF gains in non-English directions. 2) However, the performance of the boosted model is quite close to the upper bound in all of the non-English directions, which outperforms the baseline by a margin of +14.3 ChrF. 3) In the 16 English-centric directions, both the performance of the upper bound and the boosted models drop compared to the baseline, which reveals trade-off effects from English-centric to non-English directions. However, the boosted method only shows slight degradation in a few English-centric directions (e.g., en-fi and en-it in Figure 4), while the upper bound model drops in most of them. Detailed scores, including those in other metrics, are provided in Appendix A.4. In short, the boosted model achieves evident non-English gains (+14.3 ChrF on 870 language pairs) with a negligible cost in English-centric directions (-0.3 ChrF).
To further investigate the scale effect of fine-tuning data, we conduct another experiment on Europarl-8. Note that because Europarl-8 is fully multi-parallel, we can easily build the corresponding fine-tuning data covering full directions at different scales. Here, we sample multiple sets of fine-tuning data with different seeds 3 times, ranging from 1 to 12.8K samples. The average results in ChrF for each scaling step are provided in Figure 5. Also, we report the corresponding off-target ratio evaluated by fastText [^6] https://github.com/facebookresearch/fastText (Joulin et al., 2017).
The starting point (fine-tuning with 0 samples) denotes the original English-centric model trained on Europarl-8. Notably, a high off-target ratio (51.8%) exists at this point. Intriguingly, even with just 1 multi-parallel sample for fine-tuning, highly evident zero-shot improvements can be found (from 25.7 to 36.2 ChrF). Meanwhile, the off-target issue is nearly addressed (from 51.8% to 1.9%). From 1 to 100 samples, we can still observe evident zero-shot capability boosting (from 36.2 to 39.3 ChrF), while the off-target change is marginal. Further scaling up the dataset from the point of 100 samples shows nearly linear performance gains.
Figure 5: The scaling effects of the fine-tuning data. We show both the zero-shot performance and the off-target ratio for each scaling step.
4 Analysis
@@ 4.1 Does Multi-Parallelism Matter?
In Section 3, we showed the surprising boosting effects brought by the additional tiny multi-parallel samples. However, does the semantically equivalent signal matter here? In this section, we further explore whether the use of multi-parallel data, as opposed to trivial pairwise data, for fine-tuning is crucial for enhancing performance.
We use a subset of 5 languages (20 directions) to fine-tune the English-centric model built on EC30. The fine-tuning data is also built by NTREX, where we randomly map the 1997 multi-parallel samples into 10 buckets (100 samples for each). Then, we construct pairwise data in the following two ways:
@@
    Multi-Parallel: We constructed Pairwise samples in 20 directions using the multi-parallel data within one certain bucket.
@@
    Multi-Directional: For each bucket, we construct fine-tuning samples for a specific language pair only (2 directions), also resulting in 20 directional data.
Note that the size of the bitext in settings (a) and (b) are identical. However, in (a), each sentence has semantically equivalent counterparts in all other languages. In contrast, in (b), each sentence has only one unique counterpart, resulting in trivial pairwise data, i.e., multi-directional data.
To cover different language families and resource levels, we choose 5 languages, i.e., DE, FR, RU, HE, and AR. In Table 3, we show the results that fine-tuning the EC30-based English-centric model with data in (a) multi-parallel and (b) multi-directional settings, respectively. Firstly, compared to the baseline model, evident improvements can be observed for non-English translation in both settings. Furthermore, for all the groups, fine-tuning in setting (a) lags behind that in setting (b). It reveals that the boosting effects are not inherently reliant on multi-way semantic equivalence, showing substitutability when no fully multi-parallel samples exist. Especially, for the low-resource-related group, we observe 1.0+ ChrF drops. We attribute this point to the potential semantic-level redundancy in multi-parallel data in the view of multilingualism.
@@ 4.2 The Role of Semantic and Syntactic Information in Fine-tuning Data
Considering that a small amount of fine-tuning data, e.g., 100 or even one single sample, can still evidently enhance overall zero-shot performance, a related question arises: To what extent do these improvements stem from the intrinsic information inherent in the data itself? In this section, we provide some insights into the role that the semantics and syntactic of fine-tuning data play in the unexpected enhancements for zero-shot translation.
@@
For each pair in all directions, we sample uniformly from 1 to 1000 multiple times and concatenate them to a certain length, then copy it for the source and target side as a number translation sample. Given no semantic other than numerical information is contained in this setting, we try to check whether the enhancements stem from the information other than the data itself, e.g., the tags.
@@
We utilize bilingual dictionaries from MUSE  [^7] https://github.com/facebookresearch/MUSE (Lample et al., 2018) to build word pairs for all of the directions. MUSE contains 110 English-centric bilingual dictionaries and all languages of Europarl-8 are included. We first select the intersection of English words for the 7 involved English-centric dictionaries. Then, we extend them by mapping words paired with the same English words together [^8] Specifically, for each one-to-many mapping that exists, we randomly select a one-to-one mapping.. E.g., given an EN-DE pair {bike, Fahrrad} and an EN-NL pair {bike, fiets}, we can build a new DE-NL word pair {Fahrrad, fiets}. Finally, we built 28 dictionaries (16,737 word pairs for each) covering all 56 directions.
@@
We use 100 randomly sampled multi-parallel instances from NTREX to construct pairwise data covering all directions.
To ensure a fair comparison, we maintain similar surface information across the three datasets, such as aligning the number of tokens in the number-pair and word-pair datasets with the English portion in the sentence-pair dataset. Table 4 shows the corresponding fine-tuning results: 1) Fine-tuning with number-pair data results in marginal improvements. Conversely, fine-tuning with word-pair data leads to noticeable zero-shot enhancements (+10.1 ChrF). Simultaneously, the off-target ratio also decreases to an acceptable level. It means that semantic information, particularly at the lexical level, plays an important role here. 2) When using sentence-pair data (NTREX) to fine-tune, considerable further improvements compared to word-pair counterparts can be observed, showing that syntactic-level information also matters here.
@@ 5 Conclusion
In this paper, we show that the zero-shot performance of an English-centric MMT model can be easily boosted by tiny multi-parallel data, e.g., on EC30, +21.7 ChrF average gains can be achieved by fine-tuning using 100 samples from NTREX, meanwhile preserving the English-centric performance (see Section 3.3). More surprisingly, we show that fine-tuning in a small portion (10%) of directions can achieve comparable improvements to fully fine-tuning (see Section 3.4), which are even close to the upper-bound model, i.e., complete translation (see Section 3.5).
In terms of using language tags, previous research (Wu et al., 2021) shows that the two-tag model lags behind the one-tag counterpart evidently for zero-shot capability. However, we show that fine-tuning can easily address the former’s problem. Moreover, the final performance of the two-tag model consistently surpasses the one-tag model in both zero-shot and English-centric directions across multiple metrics (see Section 3.3).
We also question early works (Zhang et al., 2020; Yang et al., 2021; Sennrich et al., 2023; Chen et al., 2023) that consider the off-target issue as a challenging problem for MMT. This paper shows that the off-target issue can be easily addressed by fine-tuning with tiny (even single) multi-parallel samples. Lastly, we shed some light on the impact of different types of information from the fine-tuning data on the final performance.
Given the clear advantages of the proposed method, we encourage the community 1) to consider the use of the fine-tuning method as a strong baseline for zero-shot translation in the future, especially for the two-tag setting, and 2) to construct more comprehensive and high-quality multi-parallel datasets that cover real-world demands.
@@ A.2 Detailed Results on EC30
A.3 Detailed Results: More Data or More Directions?
@@
A.4 Detailed Results: How Close to the Upper Bound?
@@
Table 12: Zero-shot performance (SacreBLEU) on the EC30 dataset (61M sentence pairs) in two tagging fashions, grouped by High-, Medium, and Low-resource, respectively. $\Delta$-100 and $\Delta$-All mean the corresponding performance changes compared with the baselines.
@@
Table 14: Zero-shot performance (COMET) on the EC30 dataset (61M sentence pairs) in two tagging fashions, grouped by High-, Medium, and Low-resource, respectively. $\Delta$-100 and $\Delta$-All mean the corresponding performance changes compared with the baselines.

### 26 Feb 2024 v2

@@ 3.5 More Data or More Directions?
In Section 3.4, we showed that fine-tuning an English-centric model with a small amount of bitext derived from NTREX (covering all directions) yields substantial zero-shot improvements. A natural assumption is that the improvement in each direction is triggered by the corresponding directional data. In this section, we investigate whether this is true, i.e., what will happen if we only cover a subset of translation directions during fine-tuning.
We conduct experiments on the same English-centric model trained on EC30, see Section 3.4, and control the scale of the fine-tuning data in the following two settings: (a) We conducted a random sampling of 100 multi-parallel NTREX sentences to construct pairwise data to cover all 870 directions [^5] These directions include English-centric ones while excluding Occitan-related ones as it is not available in NTREX.. Then, we randomly sampled {10%, 20%, 40%, 80%} directions for fine-tuning. (b) We fixed the 10% of directions as mentioned in (a) and conducted a random sampling of {100, 200, 400, 800} multi-parallel NTREX instances to construct the fine-tuning set for the corresponding directions.
@@
Figure 3: ChrF improvements of the upper bound and boosted models over the English-centric baseline on the Europarl-8 dataset. It is clear that the overall non-English capability of the boosted model is close to the upper bound (complete translation), meanwhile, it also holds the performance in English-centric directions.
We show all of the corresponding fine-tuning results in Figure 2. Surprisingly, when fixing the size of the multi-parallel samples to 100 and then increasing the fine-tuning directions from 10% to 80%, no improvement is observed for any resource group (Figure 2-a). Fine-tuning in randomly sampled 10% directions using 100 samples achieves comparable overall results to fine-tuning in all directions (Boost-100). However, when we fix the directions to 10% and increase the multi-parallel sample size from 100 to 800, consistent improvements for all groups can be observed (Figure 2-b). This shows that the overall improvements are not sensitive to the number of directions, at least when the directions extend to a certain scale, like 10%.
In Appendix A.4.1, we further show that when we limited the fine-tuning direction set to fall in a specific family (Germanic), overall improvements just slightly lag behind that of fully fine-tuning, showing surprising cross-lingual transfer ability.
3.6 How Close to the Upper Bound?
In this section, we show to what extent our fine-tuning method can approximate an upper bound. Here, we consider the performance of complete translation, i.e., training with fully multi-parallel data, as the "upper bound", since identical scales of non-English bitext cover all of the directions that the English-centric counterparts can not cover.
We conduct experiments on Europarl-8, see Section 3.2, where 8-way aligned data are available. Both English-centric and complete translation models are trained based on it. Note that we reuse the former’s vocabulary for the latter to ensure a fair comparison. Also, we present the results after fine-tuning the English-centric model using full-direction pairwise data constructed from NTREX.
In Figure 3, we show that 1) the upper-bound performance, i.e., that of the complete model, surpasses the baseline by a large margin, resulting in +17.4 average ChrF gains for non-English directions. 2) However, the boosted model’s performance closely approaches the upper bound in all non-English directions. 3) For the 14 English-centric directions, both the upper bound and the boosted models exhibit degradation compared to the baseline, which reveals trade-off effects from English-centric to non-English directions. However, the boosted method only slightly degrades for a few English-centric directions (e.g., en-fi and en-it in Figure 3), whereas the upper bound model drops for most. Detailed scores, including those in other metrics, are provided in Appendix A.5. In short, the boosted model achieves strong non-English gains (+14.3 ChrF) with a negligible cost in English-centric directions (-0.3 ChrF).
Figure 4: Zero-shot performance and off-target ratio on Europarl-8 at each scaling step. The green solid line denotes the quality improvements of the translation samples that have no off-target issue.
4 Analysis
4.1 Off-Target and Fine-Tuning Data Size
The off-target problem is often viewed as a primary cause that impairs the zero-shot capability (Zhang et al., 2020; Yang et al., 2021; Sennrich et al., 2023; Chen et al., 2023). In this section, we delve into the impact of fine-tuning data size on off-target ratios and final performance. Moreover, we disentangle the gains of the already on-target translations, showing the extent to which the enhancements are beyond alleviating the off-target issue. Note that since Europarl-8 is fully multi-parallel, we can readily build the corresponding full-direction fine-tuning data at different scales.
@@ 4.1 Off-Target and Fine-Tuning Data Size
Here, we sample multiple sets of multi-parallel instances from the training set of Europarl-8, ranging from 1 to 12.8K with different seeds 3 times. The average results in ChrF after fine-tuning for each scaling step are provided in Figure 4. Also, we report the corresponding off-target ratio evaluated by fastText [^6] https://github.com/facebookresearch/fastText (Joulin et al., 2017) following previous works (Yang et al., 2021; Costa-jussà et al., 2022).
The blue solid line shows the overall zero-shot performance at each scaling step, where the starting point (fine-tuning with 0 samples) denotes the performance of the original English-centric model. Notably, a high off-target ratio (51.8%) exists at this point. Surprisingly, even fine-tuning with just one multi-parallel sample, very strong overall zero-shot improvements can be obtained (from 25.7 to 36.2 ChrF). Meanwhile, the off-target issue is almost completely resolved, dropping from 51.8% to 1.9%. Increasing from 1 to 100 samples, we can still observe clear zero-shot capability boosting (from 36.2 to 39.3 ChrF), while the off-target change is marginal. Further scaling up fine-tuning data from the point of 100 samples shows nearly linear performance gains.
We further disentangle the evaluation set into on-target parts [^7] Note that the on-target sample size varies across directions, with an average of 488 samples per direction. for each language direction, where each source sample is already translated into the correct direction when evaluating the English-centric model. In Figure 4, the green solid line denotes the average performance on the on-target part. It is easy to see that the improvements brought from fine-tuning with one single sample are still strong, even after isolating the impact of off-target issues. As the number of fine-tuning samples increases, on-target improvements closely follow the trend of the overall improvements, further showing that the overall improvements are not only due to resolving the off-target issue.
@@ 4.2 Does Multi-Parallelism Matter?
We have shown remarkable boosting effects obtained from adding just a tiny amount of multi-parallel data. But, does the data have to be multi-parallel? In this section, we explore whether utilizing multi-parallel data, instead of just pairwise data, for fine-tuning is vital for significant enhancements.
To this end, we fine-tune the English-centric model built on EC30 in 5 languages (20 directions). The fine-tuning data is also built from NTREX. However, we control the resulting bitext distribution. Firstly, we randomly map the 1,997 multi-parallel samples into 10 buckets (roughly 100 samples for each). Then, we construct pairwise data in the following two ways:
@@
    Multi-Parallel: We constructed pairwise samples in the 20 directions using the multi-parallel data in one randomly picked bucket.
@@
    Multi-Directional: For each bucket, we construct fine-tuning samples for a specific language pair only (2 directions), also resulting in the 20 translation directions.
Note that the size of the bitext in settings (a) and (b) are identical. In (a), each sentence has semantically equivalent counterparts in all other languages. However, in (b), each sentence has only one counterpart, resulting in simple pairwise data.
To cover different language families and resource levels, we choose DE, FR, RU, HE, and AR as these 5 languages. In Table 3, we show the results that fine-tune the EC30-based English-centric model with data in (a) multi-parallel and (b) multi-directional settings, respectively. Firstly, compared to the baseline model, clear improvements can be observed for zero-shot translation in both settings. Meanwhile, in all groups, the performance in setting (a) closely trails but never surpasses that in setting (b). It shows that the boosting effects do not depend on multi-way semantic equivalence, showing simple multi-directional data is sufficient in case fully multi-parallel samples do not exist.
@@ 4.3 The Role of Semantic and Syntactic Information in Fine-Tuning Data
Considering that a small amount of fine-tuning data, e.g., 100 or even one single sample, can still substantially enhance overall zero-shot performance, a related question arises: To what extent do these improvements stem from the intrinsic information inherent in the data itself? In this section, we provide some insights into the role that the semantics and syntactic of fine-tuning data play in the unexpected improvements for zero-shot translation.
@@
For each direction, we perform uniform sampling of digits (ranging from 1 to 1000) multiple times, concatenating them to a certain length. Then, it is replicated for both the source and target sides, forming a number translation sample, as shown in Figure 6. Given no semantic other than numerical information is contained in this setting, we try to check whether the improvements stem from factors other than the data itself, e.g., the tags.
@@
We utilize bilingual dictionaries from MUSE  [^8] https://github.com/facebookresearch/MUSE (Lample et al., 2018) to build word pairs for all of the directions. MUSE contains 110 English-centric bilingual dictionaries and all languages of Europarl-8 are included. We first select the intersection of English words for the 7 involved English-centric dictionaries. Then, we extend them by mapping words paired with the same English words together. [^9] Specifically, for each one-to-many mapping that exists, we randomly select a one-to-one mapping. E.g., given an EN-DE pair {bike, Fahrrad} and an EN-NL pair {bike, fiets}, we can build a new DE-NL word pair {Fahrrad, fiets}. Finally, we built 28 dictionaries (16,737 word pairs for each) covering all 56 directions.
@@
We use 100 randomly selected multi-parallel samples from NTREX to construct pairwise data covering all directions.
To ensure a fair comparison, we maintain similar surface information across the three datasets, such as aligning the number of tokens in the number-pair and word-pair datasets with the English portion in the sentence-pair dataset. Table 4 shows the corresponding fine-tuning results: 1) Fine-tuning with number pairs results in marginal improvements. Conversely, fine-tuning with word pairs leads to noticeable zero-shot improvements (+10.1 ChrF). Simultaneously, the off-target ratio also decreases to an acceptable level. This means that semantic information, particularly at the lexical level, plays an important role here. 2) When using sentence-pair data (NTREX) to fine-tune, considerable further improvements compared to word-pair counterparts can be observed, showing that syntactic-level information also matters.
@@ 5 Conclusion
In this paper, we show that the zero-shot performance of an English-centric MMT model can be easily boosted by a tiny amount of multi-parallel data. On EC30, +21.7 ChrF average gains can be achieved by fine-tuning using 100 samples from NTREX, meanwhile preserving the English-centric performance, see Section 3.4. More surprisingly, we show that fine-tuning on a small portion (10%) of directions can achieve comparable improvements to full-direction fine-tuning, see Section 3.5, which are even close to the ideal but impractical upper-bound model, see Section 3.6.
In terms of using language tags, we show that fine-tuning can address the two-tag model’s performance degradation in zero-shot directions (Wu et al., 2021). Moreover, the final performance substantially surpasses that of the one-tag model across multiple metrics, see Section 3.4.
We also question earlier findings (Zhang et al., 2020; Yang et al., 2021; Sennrich et al., 2023; Chen et al., 2023) that consider the off-target issue as a challenging problem for MMT. This paper shows that the off-target issue can be easily addressed by fine-tuning with tiny (even single) multi-parallel samples. Lastly, we shed some light on the impact of different types of fine-tuning data on the final performance.
Given the clear advantages of our proposed method, we encourage the community 1) to consider the use of fine-tuning as a strong baseline for zero-shot translation in the future, especially for the two-tag setting, and 2) to construct more comprehensive and high-quality multi-parallel datasets that cover real-world demands.
Limitations
Multi-parallel data are normally built in a way that translates the same English data into multiple other languages by professional human translators. Hence, in the resulting non-English fine-tuning data, both the source and target side are translated instead of using the original text. This may exacerbate potential drawbacks in certain directions, such as translations
Broader Impact
MMT systems have significant progress recently. However, potential challenges such as mistranslation or off-target issues still exist. Moreover, the fairness problem also arises, e.g., the generation ability is not guaranteed to be fair across languages or demographic features, which may run the risk of reinforcing societal biases, e.g., race bias.
@@ A.1 Detailed Dataset Description
A.2 Training Setting
For all of the English-centric training, the learning rate is 5e-4 with 4,000 warmup steps and a inverse sqrt decay schedule. All dropout rates and label smoothing are set to 0.1. In the case of EC30 and Europarl-8, the batch size is set as 8,196 tokens, accumulating gradients 20 and 8 times, respectively. Also, data from different language pairs are sampled with a temperature of 5.0 and 2.0, respectively. The same temperature is applied to both BPE building and MMT training periods. We train all models with an early-stopping strategy [^10] Patience is set to {10, 20}, i.e., training stops if performance on the validation set does not improve for the last {10, 20} checkpoints, with 1,000 steps between checkpoints. and evaluate by using the best checkpoint as selected based on the loss on the development set.
For fine-tuning, all parameters are kept the same as those in training, except for 1) we set batch accumulation as 1 in Section 4 as extremely small fine-tuning data is used, and 2) we set patience as 3 for quick experiments.
Note that we use 4 A6000 GPU cards for English-centric training with FP16 optimization, which means the actual batch size is also 4 times bigger. For fine-tuning, we use a single A6000 GPU card.
@@ A.3 Detailed Results on EC30
A.4 More Data or More Directions?
A.4.1 Limited Fine-tuning Direction Set
A.4.2 Detailed Results: More Data or More Directions?
@@
A.5 Detailed Results: How Close to the Upper Bound?
@@
A.6 Number Pairs
The synthetic number pairs and word pairs are illustrated in Figure 6.
@@ A.6 Number Pairs
Table 12: Zero-shot performance (SacreBLEU) on the EC30 dataset (61M sentence pairs) with two language tag strategies, grouped by High-, Medium, and Low-resource, respectively. $\Delta$-100 and $\Delta$-All mean the corresponding performance changes compared with the baselines.
@@
Table 14: Zero-shot performance (COMET) on the EC30 dataset (61M sentence pairs) with two language tag strategies, grouped by High-, Medium, and Low-resource, respectively. $\Delta$-100 and $\Delta$-All mean the corresponding performance changes compared with the baselines.
@@
Source                                                       Target 961 271 4 137 146 37 124 498 28 73                   961 271 4 137 146 37 124 498 28 73 323 135 42 121 324 499 17 11 37 33                   323 135 42 121 324 499 17 11 37 33 713 93 331 225 51 7 375 223 12 192                   713 93 331 225 51 7 375 223 12 192 …                                                              …
Figure 6: Illustration of number pairs.